The combustion of 1.50g sample of the rocket fuel hydrazine N2H4 released 30.4kj of heat.calculate the molar heat of combustion of hydrazine


1
Expert's answer
2021-11-05T09:31:53-0400


Specific heat = 5680J/°C

q= 5680×30.4

= 172672J= 172.672kJ

1.50/32= 0.04688molea of hydrazine

molar heat of combustion of hydrazine= 172.672/0.04688

= 3683.28kJ/mole